# ARC: Argument Representation and Coverage Analysis for Zero-Shot Long Document Summarization with Instruction Following LLMs

### v1

@@ Abstract
Integrating structured information has long improved the quality of abstractive summarization, particularly in retaining salient content. In this work, we focus on a specific form of structure: argument roles, which are crucial for summarizing documents in high-stakes domains such as law. We investigate whether instruction-tuned large language models (LLMs) adequately preserve this information. To this end, we introduce Argument Representation Coverage (ARC), a framework for measuring how well LLM-generated summaries capture salient arguments. Using ARC, we analyze summaries produced by three open-weight LLMs in two domains where argument roles are central: long legal opinions and scientific articles. Our results show that while LLMs cover salient argument roles to some extent, critical information is often omitted in generated summaries, particularly when arguments are sparsely distributed throughout the input. Further, we use ARC to uncover behavioral patterns—specifically, how the positional bias of LLM context windows and role-specific preferences impact the coverage of key arguments in generated summaries, emphasizing the need for more argument-aware summarization strategies.
@@ 1 Introduction
LLMs have made remarkable progress in text summarization, often generating summaries preferred by human evaluators in domains like news Zhang et al. (2024); Liu et al. (2024b). However, summarization in structured, highly informative domains presents unique challenges that remain underexplored. One prominent challenge is preserving salient argument roles in the generated summaries, which can be challenging as arguments can be sparsely distributed across the input Elaraby and Litman (2022). The ability of LLMs to selectively retain argument roles is therefore a crucial test of their utility in generating reliable summaries in high-stakes domains.
Prior work in legal summarization has shown that explicitly modeling argument roles—either during finetuning Fabbri et al. (2021a); Elaraby and Litman (2022) or through post hoc re-ranking Elaraby et al. (2023)—improves the coverage of critical argumentative content. These findings suggest that pretrained language models (PLMs) may struggle to capture structured discourse elements such as arguments without targeted supervision. However, it remains an open question whether instruction-tuned LLMs, trained with broad and often general-purpose supervision, can inherently identify and preserve salient argumentative information without explicit signals about saliency or additional tuning. In this work, we go beyond conventional summary evaluation metrics such as fluency, factuality, and coherence to address a core question: Do LLMs effectively prioritize and preserve the most salient argumentative content in their summaries?
To address this question, we introduce Argument Representation Coverage (ARC), a framework for evaluating how well LLM-generated summaries capture salient arguments. ARC measures coverage at three levels of granularity: (1) Argument Set Coverage (), assessing collective coverage of full salient argument set; (2) Independent Role Coverage (), evaluating each argument role separately as atomic units; and (3) Subatomic Coverage (), examining fine-grained factual units within roles.
Figure 1 presents an example of how ARC operates across multiple levels of abstraction. While provides a single holistic score capturing overall argument coverage, it fails to offer fine-grained insight into which specific arguments are preserved or omitted—information critical for downstream analyses such as bias detection. addresses this by assessing the preservation of each argument role independently, although it considers an argument entirely unsupported if even one element is missing. To mitigate this limitation, decomposes arguments into subatomic factual units, enabling a more continuous and informative scoring mechanism. This multi-level approach is designed to overcome the limitations of coarse Likert-scale ratings, which have been shown to suffer from low inter-annotator agreement in prior work Elaraby et al. (2023); Krishna et al. (2023a).
@@
Prior work highlights that LLMs often exhibit positional biases—favoring content from the beginning or end of long documents Liu et al. (2024a); Ravaut et al. (2024a); Wan et al. (2024). Furthermore, it remains unclear whether LLMs also favor certain types of salient information that share the same structure (e.g., argument roles) during summarization. Using ARC, we conduct two key analyses: (1) How does the position of arguments in the source document affect their inclusion in summaries? and (2) Are certain argument roles disproportionately favored over others? The latter analysis is crucial, as uneven role coverage may lead to biased or incomplete summaries that misrepresent the original discourse. We conduct our experiments across two domains where argument structure plays a central role in understanding the essence of the document: Long Legal Opinions and Scientific Articles.
Our contributions are in two folds: (1) we propose ARC, a multi-granularity evaluation framework for evaluating argument coverage; and (2) we present a systematic analysis of how positional and role-specific biases affect LLMs’ ability to preserve salient argumentative content in summaries.
@@ 2 Related Work
Content selection remains a core challenge in summarization. Trienes et al. (2025) found weak alignment between LLMs’ saliency preferences and human judgments. While LLMs can produce summaries preferred over human references in domains like news Zhang et al. (2024); Liu et al. (2024b), they still benefit from content planning. For example, Adams et al. (2023) showed that planning entity mentions improves the information density in GPT-4 generated summaries at the same summary length when compared to summaries generated without entitiy planning. We extend this line of work by treating argument roles as a structured form of saliency and analyzing their preservation in LLM-generated summaries.
Limitations of LLMs in Long-Document Summarization.
@@
LLMs face persistent issues when summarizing long texts, notably the U-shaped positional bias—favoring content at the beginning and end while neglecting the middle Ravaut et al. (2024b). This leads to degraded faithfulness in long-form outputs Wan et al. (2024). We expand this analysis by quantifying how positional bias affects the coverage of salient argumentative content.
Argument Mining and Abstractive Summarization.
@@
Incorporating argument structures into summarization has shown promise across domains, including dialogues Fabbri et al. (2021a), legal texts Xu et al. (2020, 2021); Elaraby and Litman (2022); Elaraby et al. (2023) and scientific documents Fisas et al. (2016). We build on this by assessing whether instruction-tuned LLMs can cover salient arguments without the external argument role information, particularly in summarizing legal and scientific texts.
Evaluation Metrics for Long-Form Summarization.
Standard metrics often fall short in reflecting human preferences, especially for long documents Fabbri et al. (2021b); Krishna et al. (2023b). To improve reliability, recent work has introduced unit-based metrics—such as Atomic Content Units (ACUs) Krishna et al. (2023a) and structured factuality scores Min et al. (2023); Yang et al. (2024)—to reduce subjectivity. Extending this idea, we propose ARC, which uses argumentative structures as evaluation units and introduces subatomic granularity to assess fine-grained argument coverage.
3 Datasets111More analysis and examples in Appendix A.
We employ two datasets that include both argument role annotations and reference summaries: CANLII Xu et al. (2021), representing the legal domain, and DR. INVENTOR (DRI) Fisas et al. (2016), representing the scientific domain. An overview of dataset statistics is presented in Table 1. Both datasets consist of long-form documents paired with long-form reference summaries that average words Krishna et al. (2023b).
| Dataset | Argument Role | Example |
@@
| CANLII | Issue | Damage to both vehicles exceeded the insurance deductibles and both parties claim damages against each other. |
| Conclusion | Fault for this accident was attributed to the defendant and to the plaintiff. | |
| Reason | Jurisdictional error is not to be equated with error of law. | |
@@
| DRI | Own Claim | Semi-Lagrangian contouring offers an elegant and effective means for surface tracking with advantages over competing methods. |
@@
| Background Claim | Accurate modeling of human motion remains a challenging task. | |
| Data | Animation is constrained due to hardware constraints. |
@@ 3.2 Scientific Articles: DRI
The DRI dataset consists of computer graphics articles, each annotated at the sentence level for rhetorical roles and paired with three human-written summaries. Notably, these rhetorical roles are not necessarily argumentative, making it challenging to assess argument coverage. To address this limitation, the extended version of the dataset, SCI-ARG Lauscher et al. (2018), enriches the DRI annotations by incorporating argument role annotations and their relations. These annotations follow a modification of the -argument roles described in Toulmin model Toulmin (2003), by reducing them into three types: Own Claim: Sentences that directly support the author’s central argument. Background Claim: Sentences that reference prior research or established domain knowledge. Data: Empirical evidence that supports or refutes claims, such as experimental results or literature citations. An example of each role is shown in Table 2.
@@ 4.1 Overview and Notations
Given a generated summary and a set of salient arguments , we define a coverage function , where: . Here, represents the evaluation unit against , where: . represents the full set of salient arguments in the document. is an individual argument . is a subatomic factual unit, where and is the set of atomic facts derived from argument .
Full Argument Set Coverage
Following Elaraby et al. (2024), we define based on a Likert-scale annotation (1-4), which is then normalized to a range:
where is the Likert score assigned to the argument set coverage.
Independent Argument Role Coverage
Each argument is evaluated independently as:
Subatomic Argument Coverage
An argument is further decomposed into subatomic fact units , where for :
The scores obtained by at different granular levels provide a deeper assessment of model behavior on the argument level. Table 3 shows ARC scores computation for an number of salient arguments.
| Granularity | Computation |
| Full Set | |
| Individual Roles | |
| Subatomic Units |
| Metric | Metric Type | Expert 1 | Expert 2 | Average | |||
| ROUGE-1 | Lexical | 0.3455 | 0.4809 | 0.3382 | 0.4495 | 0.3757 | 0.5318 |
| ROUGE-2 | 0.3233 | 0.4150 | 0.2860 | 0.4132 | 0.3291 | 0.4734 | |
| ROUGE-L | 0.2824 | 0.3874 | 0.3463 | 0.4455 | 0.3418 | 0.4764 | |
| BERTScore | Semantic | 0.3187 | 0.4092 | 0.2921 | 0.3977 | 0.3344 | 0.4756 |
| (sent) | Entailment | 0.3676 | 0.5157 | 0.3567 | 0.4368 | 0.3204 | 0.4654 |
| (doc) | 0.4512 | 0.4894 | 0.2617 | 0.3438 | 0.3747 | 0.4758 | |
| (sent) | 0.3204 | 0.4654 | 0.3567 | 0.4368 | 0.3676 | 0.5157 | |
| (doc) | 0.4512 | 0.4894 | 0.2617 | 0.3438 | 0.3747 | 0.4758 | |
| LLM-Judge | 0.6713 | 0.7288 | 0.4072 | 0.4793 | 0.5867 | 0.6898 | |
| 0.4884 | 0.6034 | 0.3453 | 0.4527 | 0.4474 | 0.5971 | ||
| 0.5806 | 0.7023 | 0.5135 | 0.6353 | 0.6025 | 0.7560 | ||
| 0.5025 | 0.6026 | 0.4593 | 0.5727 | 0.5142 | 0.6642 | ||
| 0.5213 | 0.6507 | 0.4347 | 0.5304 | 0.5202 | 0.6959 | ||
4.2 Choice of
Given the limited availability of large, diverse argument coverage datasets, we use the annotated dataset from Elaraby et al. (2024), comprising legal opinions from the CANLII dataset with corresponding summaries. Each summary is annotated for argument coverage using a 4-point Likert scale at the full argument set level.333See Appendix B for scale definitions.
Computing using LLM-judge
We use GPT-4o with temperature sampling as an automated evaluator, following prior work that demonstrated its strong correlation with human judgments in summarization tasks Liu et al. (2023b).
: We prompt the model to assign a Likert score (1-4) based on coverage guidelines from Elaraby et al. (2024). Scores are then normalized to for comparability with other metrics.
: The model assigns binary scores (1 = fully supported, 0 = missing or inconsistent) for each argument .
: Building on recent factuality frameworks that decompose summaries into atomic fact units for grounded evaluation Min et al. (2023); Lee et al. (2024); Tang et al. (2024); Yang et al. (2024), ARC instead decomposes key arguments need to be retained. This shift focuses evaluation on the completeness of argument roles rather than factuality evaluation. Unlike prior completeness work, which treats all information uniformly, argument roles impose a structured prioritization, which is sparse in legal cases and dense in scientific articles, enabling more targeted analysis to indicate if it’s still worthy including structured information such as argument roles for a more complete summary.
Following the decomposition method of Yang et al. (2024), each argument is decomposed into a set of factual units using GPT-4o. Units are filtered by a fine-tuned entailment model (DeBERTa; He et al. )444https://huggingface.co/cross-encoder/nli-deberta-v3-base to keep only entailed facts555Decomposition details in Appendix C.. Each unit is then verified against the summary , yielding , where indicates support and specifies the error type (missing or non-factual). This enables fine-grained argument coverage evaluation while distinguishing hallucination from information loss666Evaluation prompts in Appendix D..
Reducing LLM-Judge Cost Evaluating argument coverage at scale using GPT-4o is costly, particularly for the legal dataset (CANLII), where each document contains numerous argument units and their subatomic facts. To mitigate this cost, we train classifiers to approximate GPT-4o judgments, thereby reducing reliance on LLM-based evaluation.
We sample cases from CANLII and use GPT-4o to compute and for all models. This data serves as the training set for two classifiers. The first classifier, , is a binary classifier that predicts argument-level support, trained to approximate the judgments of GPT-4o such that . The second classifier, , is a three-way classifier that predicts whether a subatomic fact is supported, missing, or non-factual, ensuring that .
We explored several models, including DeBERTa (base and large), LegalBERT Chalkidis et al. (2020), and BigBirdRoBERTa Zaheer et al. (2021). Among all models, the DeBERTa-large demonstrated the highest performance, achieving an F1-macro score of for subatomic facts evaluation and for role evaluation against summary predictions, using -fold cross-validation 777Training setup and evaluation against GPT-4o are provided in Appendix E..
ARC Scores Against Human Evaluation We compare ARC scores with lexical and semantic metrics from Elaraby et al. (2024), as well as entailment-based metrics, particularly SummaC Laban et al. (2022) (both zero-shot and convolution-based), which measures alignment between source articles and summaries. For entailment evaluation, argument roles are treated as the hypothesis and generated summaries as the premise, with SummaC computed at both document- and sentence-level granularities. Table 4 shows that ARC consistently outperforms lexical, semantic, and entailment metrics in correlating with expert annotations, with one exception: for against Expert 2. The lower performance of suggests that treating a full argument as an atomic unit introduces excessive penalization for minor omissions and inconsistencies, leading to misalignment with holistic human judgments. In contrast, achieves the highest correlations with Expert 2 and the overall expert average, enabling getting a nuanced and interpretable score without losing the holistic overall coverage.
When using trained classifiers, we observe a drop in performance but an improvement in , which is likely due to distributional shifts between training and human-evaluated summaries. Nonetheless, the stricter nature of appears to amplify small deviations, reducing alignment with expert assessments. Given its strong expert correlation and interpretable error predictions, we adopt ARC as the primary metric for coverage evaluation, providing a reliable and fine-grained lens on model performance.

### v2

@@ Abstract
We introduce Argument Representation Coverage (ARC), a bottom-up evaluation framework that assesses how well summaries preserve salient arguments, a crucial issue in summarizing high-stakes domains such as law. ARC provides an interpretable lens by distinguishing between different information types to be covered and by separating omissions from factual errors. Using ARC, we evaluate summaries from eight open-weight LLMs in two domains where argument roles are central: long legal opinions and scientific articles. Our results show that while LLMs capture some salient roles, they frequently omit critical information, particularly when arguments are sparsely distributed across the input. Moreover, ARC uncovers systematic patterns—showing how context window positional bias and role-specific preferences shape argument coverage—providing actionable guidance for developing more complete and reliable summarization strategies.
@@ 1 Introduction
LLMs have achieved remarkable progress in abstractive summarization, producing summaries that are fluent, coherent, and often preferred by human evaluators in domains such as news Zhang et al. (2024); Liu et al. (2024b). This progress has shifted the focus of summarization research from how to generate fluent summaries to how to properly evaluate them. Evaluation is especially critical because, despite their fluency, LLMs frequently hallucinate or omit key content Huang et al. (2025); Kim et al. , undermining their reliability in high-stakes domains.
In this work, we focus on the evaluation of long-context summaries generated by instruction-following LLMs, focusing on whether they preserve salient argumentative content. In crucial domains such as law, the most essential information is conveyed through argument roles (the components of reasoning and main claims in the text). Preserving these roles is particularly challenging in long documents, where arguments are sparsely distributed across thousands of tokens Elaraby and Litman (2022). The ability of LLMs to identify and retain argument roles is therefore a crucial test of their utility for reliable summarization in high-stakes settings.
Why argument-based evaluation? Prior work has shown that explicitly modeling argument roles can improve summarization quality Fabbri et al. (2021b); Elaraby and Litman (2022); Elaraby et al. (2023). Building on this, argument-based evaluation allows us to address a fundamental research question: RQ1: Can instruction-following LLMs inherently identify and preserve argument roles without explicit supervision? Additonally, we see two key benefits of argument-based evaluation where existing metrics fall short: interpreting which types of information are missing or misrepresented in generated summaries, and guiding future alignment strategies for domain-specific summarization through a reproducible methodology. In this work, we focus on the first point by developing argument-based evaluation methods that make coverage interpretable.
We introduce Argument Representation Coverage (ARC), 111The full codebase, along with instructions for obtaining the data, is publicly available at https://github.com/EngSalem/ARCScore.git. a framework for evaluating how well LLM-generated summaries capture salient arguments. ARC leverages atomic fact decomposition of argument roles to measure coverage hierarchically. As illustrated in Figure 1 (decomposing roles to facts), all salient arguments are first decomposed into atomic facts to assess fact-level coverage. Each atomic fact is then compared against the generated summary and labeled as either supported, missing, or contradicted. These binary fact-level judgments are aggregated to compute role-level coverage (), providing fine-grained insight into how well each argument role is preserved. Finally, role-level scores are combined to produce summary-level coverage (), a holistic measure of how effectively a summary preserves salient argument roles. We apply ARC to study coverage in two domains where argument structure is essential for understanding the core of the document: long legal opinions and scientific articles.
@@
Prior work shows that LLMs often exhibit positional biases—favoring content from the beginning or end of long documents Liu et al. (2024a); Ravaut et al. (2024); Wan et al. (2025). It is less clear, however, whether they also bias toward certain types of salient information that share the same structure (e.g., argument roles) during summarization. With ARC, we investigate additional two research questions: RQ2: How does the position of arguments in the source document affect their inclusion in summaries? and RQ3: Are certain argument roles disproportionately favored over others? To answer the latter, we leverage ARCrole to define a bias score that quantifies role-specific biases in saliency coverage.
Our analysis with ARC across eight open-weight LLMs reveals: instruction-following models still struggle with saliency coverage, particularly when argument roles are sparsely distributed, underscoring the need for further alignment; positional bias in the context window negatively impacts coverage; and bias analysis shows that LLMs exhibit systematic preferences for certain argument roles, particularly in long legal opinions.
@@ 2 Related Work
Information Saliency in LLMs’ Summaries. Content selection remains a core challenge in summarization. Trienes et al. (2025) found weak alignment between LLMs’ saliency preferences and human judgments. While LLMs can produce summaries preferred over human references in domains like news Zhang et al. (2024); Liu et al. (2024b), they still benefit from content planning. For example, Adams et al. (2023) showed that planning entity mentions improves the information density in GPT-4 generated summaries at the same summary length when compared to summaries generated without entity planning. In similar veins, Gantt et al. (2024); Walden et al. (2025) show that LLMs generate more focused and complete event-centered summaries when generation is explicitly conditioned on both the input document(s) and structured event representations. We extend this line of work by treating argument roles as a structured form of saliency and analyzing their preservation in LLM-generated summaries.
@@
LLMs in Long-Document Summarization. LLMs face persistent issues when summarizing long texts, notably the U-shaped positional bias Liu et al. (2024a)—favoring content at the beginning and end while neglecting the middle Ravaut et al. (2024). This leads to degraded faithfulness in long-form outputs Wan et al. (2025). We expand this analysis by quantifying how positional bias affects the coverage of salient argumentative content.
@@
Argument Mining in Abstractive Summarization. Incorporating argument structures into summarization has shown promise across domains, including dialogues Fabbri et al. (2021b), legal texts Xu et al. (2020, 2021); Elaraby and Litman (2022); Elaraby et al. (2023) and scientific documents Fisas et al. (2016). We build on this by assessing whether instruction-following LLMs can cover salient arguments without the external argument role information
| CANLII (Legal opinions) |
@@
| Issue: Damage to both vehicles exceeded the insurance deductibles and both parties claim damages against each other. |
| Conclusion: Fault for this accident was attributed to the defendant and to the plaintiff. |
| Reason: Jurisdictional error is not to be equated with error of law. |
| DRI (Scientific documents) |
@@
| Own Claim: Semi-Lagrangian contouring offers an elegant and effective means for surface tracking with advantages over competing methods. |
@@
| Background Claim: Accurate modeling of human motion remains a challenging task. |
| Data: Animation is constrained due to hardware constraints. |
Evaluating Long-Form Summaries. Standard metrics often fall short in reflecting human preferences, especially for long documents Fabbri et al. (2021a); Krishna et al. (2023). To improve reliability, recent work has introduced unit-based metrics—such as Atomic Content Units (ACUs) Krishna et al. (2023) and structured factuality scores Min et al. (2023); Yang et al. (2024b)—to reduce subjectivity in evaluation. Extending this idea, we propose ARC, which uses argument roles as evaluation units and introduces subatomic granularity to assess fine-grained argument coverage.
3 Datasets
We employ two datasets that include both argument role annotations and reference summaries: CANLII Xu et al. (2021), representing the legal domain, and DR. INVENTOR (DRI) Fisas et al. (2016), representing the scientific domain. An overview of dataset statistics is presented in Table 1. Both datasets consist of long-form documents paired with long-form reference summaries that average words Krishna et al. (2023). More analysis and examples are given in Appendix A.
@@ 3.2 Scientific Articles: DRI
The DRI dataset consists of computer graphics articles, each annotated at the sentence level for rhetorical roles and paired with three human-written summaries. Notably, these rhetorical roles are not necessarily argumentative. To address this limitation, the extended version of the dataset, SCI-ARG Lauscher et al. (2018), enriches the DRI annotations by incorporating argument role annotations and their relations. These annotations follow a modification of the -argument roles described in Toulmin model Toulmin (2003), by reducing them into three types: Own Claim: Sentences that directly support the author’s central argument. Background Claim: Sentences that reference prior research or established domain knowledge. Data: Empirical evidence that supports or refutes claims, such as experimental results or literature citations. An example of each role is shown in Table 2.
@@ 4.1 Overview and Notations
Decomposing Roles to Facts. For each argument role , where is the set of salient argument roles (roles that are essential to be included in the summary), we follow the decomposition algorithm of Yang et al. (2024b). Specifically, we employ a strong LLM (GPT4-o) to decompose into a set of atomic facts . These facts are further filtered using an entailment check against to eliminate over-generated facts. 222Decomposition algorithm in Appendix B.
Role-Level Coverage. Given a generated summary , we define the indicator function:
The role-level coverage for is then:
Summary-Level Coverage. The overall coverage score, ARCscore, across all roles is computed as aggregation of across :
4.2 Evaluating Factual Units ()
The main goal of the ARC framework is to provide an interpretable evaluation of coverage in generated summaries. To do so, the function distinguishes between two error types: missing facts (information omitted) and incorrectly covered facts (information misrepresented). 333Examples of atomic facts are included in Appendix C. Although both reduce completeness, separating them reveals whether errors stem from failing to prioritize salient content or from factual mistakes during generation. We implement using an LLM judge. 444See Appendix D for the evaluation prompt. We employ zero-shot LLMs, which have shown strong correlation with human judgments in summary evaluation Liu et al. (2023). Given an atomic fact and a summary, we prompt the LLM to return a pair , where indicates whether the fact is supported (covered) or not supported (missing or incorrect), and provides a categorical interpretation of the error type (no error, missing, non-factual).
4.3 Benchmarking ARCscore
To evaluate the effectiveness of ARCscore in capturing holistic summary coverage, we benchmark it against existing automatic metrics using expert-annotated data.
Expert-Annotated Data. Human-annotated datasets for fine-grained coverage evaluation are extremely scarce. To our knowledge, the only available resource is the dataset of Elaraby et al. (2024), which includes legal opinion summaries annotated for salient argument coverage on a -point Likert scale. 555See Appendix E for scale definitions. Two legal experts independently annotated the data, achieving a moderate agreement of (quadratic weighted). To improve the robustness of the annotations, we filtered out pairs with point of disagreement, resulting in article–summary pairs with improved agreement (). We use the expert average as the gold standard to compare against automatic metrics.
Automatic Metrics for Holistic Coverage. We benchmark ARC against a suite of strong automatic metrics. For all baselines, we treat the set of salient arguments as the hypothesis and the generated summary as the premise/reference. We first include the standard metrics reported in Elaraby et al. (2024): ROUGE-1, ROUGE-2, ROUGE-L, and BERTScore. Next, we add entailment-based baselines from Laban et al. (2022), where SummaC computes entailment between hypothesis and reference at both sentence-level and document-level granularity. We include both (zeroshot zs and convolution conv versions of SummaC). Finally, we incorporate an LLM-judge baseline: FactScore Min et al. (2023), which—similar to ARCscore—relies on decomposition for fact-level evaluation. We adapt FactScore for coverage by treating the generated summary as the knowledge base and the salient arguments as hypotheses to be tested. For fairness, both ARCscore and FactScore use GPT4-o as the evaluator. We report correlation with expert averages using Kendall’s and Pearson’s .
| Metric | Expert Avg. | Interpretability | ||
| Info Types | Errors | |||
| ROUGE-1 | ✗ | ✗ | ||
| ROUGE-2 | ✗ | ✗ | ||
| ROUGE-L | ✗ | ✗ | ||
| BERTScore | ✗ | ✗ | ||
| SummaCZS (sent) | ✗ | ✗ | ||
| SummaCZS (doc) | ✗ | ✗ | ||
| SummaCconv (sent) | ✗ | ✗ | ||
| SummaCconv (doc) | ✗ | ✗ | ||
| FactScore (GPT4-o) | ✗ | ✗ | ||
| ARCscore (GPT4-o) | ✓ | ✓ | ||
| Judge Model | ||
| Proprietary models | ||
| GPT-4o | 0.465 | 0.593 |
| GPT-4o-mini | 0.446 | 0.582 |
| Open-weight Instruction-following models | ||
| Qwen-2.5-7B-Instruct | 0.329 | 0.480 |
| Qwen-2.5-14B-Instruct | 0.454 | 0.601 |
| Llama-3.1-8B-Instruct | 0.463 | 0.610 |
| Mistral-8B-Instruct | 0.443 | 0.599 |
| Open-weight Reasoning models | ||
| QwQ-32B-AwQ | 0.437 | 0.596 |
| DeepSeek-R1-Distill-Qwen-14B | 0.509 | 0.638 |
Table 3 shows that among automatic metics, ARCscore achieves the highest correlation with expert judgments,666See Appendix F for correlations with each expert. outperforming standard evaluation metrics and LLM-based decomposition-based metrics such as FactScore. Crucially, ARCscore improves the holistic view of coverage over baselines, while also providing interpretability that existing metrics lack. Unlike FactScore, which provides only fact-level coverage, ARC enables multi-level analysis: role-level coverage via ARCrole, and error-type analysis distinguishing between missing and incorrect facts. This decomposition allows us to diagnose not only how much coverage is achieved, but also which roles are underrepresented and why errors occur, making ARC a more effective diagnostic tool for granular coverage evaluation.
Alternative LLM-judges. To make ARC more scalable and accessible, we explore alternatives to GPT-4o for fact-level evaluation that maintain high correlation with expert ratings while reducing inference cost and avoiding reliance on proprietary APIs that may be deprecated. We focus on two categories of open-weight models: instruction-following models. We evaluate smaller open-weight models that fit within our compute and memory budgets, including Qwen-2.5-Instruct (7B and 14B), Llama-3.1-8B-Instruct, and Mistral-8B-Instruct. Reasoning models. We further examine open-weight reasoning models, specifically QwQ-32B-AwQ and DeepSeek-R1-Distill-Qwen-14B, which have demonstrated competitive performance with proprietary instruction-following and reasoning models on complex tasks, including verification. Table 4 shows that DeepSeek-R1-Distill-Qwen-14B achieves the strongest overall correlation, surpassing even proprietary models such as GPT-4o. Among open-weight instruction-following models, Llama-3.1-8B-Instruct attains the highest correlation, outperforming GPT-4o-mini. We employ both proprietary (GPT-4o) and open-weight judge (DeepSeek-R1-Distill-Qwen-14B) to compute ARCscore. To balance cost and scalability, GPT-4o is applied to a representative subset of CANLII (CANLII100) cases and the full DRI corpus, while DeepSeek-R1-Distill-Qwen-14B is used for both CANLII100 and the full CANLII ( articles). Our goal is to ensure that the sensitivity to the underlying judge (i.e., disagreements at the fact level) does not alter the overall coverage trends observed across models.
| Model | CANLII100 | CANLII | DRI | ||
| GPT4-o | DeepSeek | DeepSeek | GPT4-o | DeepSeek | |
| Reference | 0.986 | 0.977 | 0.982 | 0.838 | 0.813 |
| Qwen-2.5-14B | 0.677 | 0.707 | 0.722 | 0.772 | 0.797 |
| Qwen-2.5-7B | 0.676 | 0.691 | 0.725 | 0.799 | 0.806 |
| Qwen-2.5-3B | 0.648 | 0.702 | 0.710 | 0.730 | 0.759 |
| Qwen-2.5-1.5B | 0.472 | 0.512 | 0.539 | 0.668 | 0.628 |
| Mistral-8B | 0.554 | 0.609 | 0.608 | 0.669 | 0.697 |
| LLaMA-3.1-8B | 0.641 | 0.642 | 0.678 | 0.749 | 0.793 |
| LLaMA-3.2-3B | 0.573 | 0.603 | 0.639 | 0.705 | 0.740 |
| LLaMA-3.2-1B | 0.451 | 0.530 | 0.545 | 0.674 | 0.631 |
5 Analyzing Summary Coverage with ARC
In this section, we use ARC to evaluate summaries produced by long-context LLMs on the CANLII and DRI datasets.